OpenStudy (anonymous):

How to find the derivative of: f(x) = x^3(x − 7)^2

OpenStudy (amistre64):

product rule is my first impression

OpenStudy (anonymous):

product rule and chain rule

OpenStudy (amistre64):

assuming its a product to begin with, and not a confuddled exponent

OpenStudy (anonymous):

how would I go about starting... im sorry, i am terrible at math haha

OpenStudy (amistre64):

\[D[fg]=f'g+fg'\]

OpenStudy (anonymous):

x^2(5x-7)(5x-7) something like that?

OpenStudy (amistre64):

..... maybe. but i cant see it as that yet

OpenStudy (anonymous):

(x-7)*

OpenStudy (anonymous):

i probably messed up, what did you get?

OpenStudy (amistre64):

f(x) = x^3 * (x − 7)^2 f'(x) = x'^3 * (x − 7)^2 + x^3 * (x − 7)'^2 f'(x) = 3x^2 * (x − 7)^2 + x^3 * 2(x − 7)(x − 7)' f'(x) = 3x^2 * (x − 7)^2 + x^3 * 2(x − 7)(1)

OpenStudy (amistre64):

just follow the definition of a product rule ....

OpenStudy (amistre64):

and notice the chain rule pops out that (x-7)'

OpenStudy (anonymous):

yeah alright let me see